OpenStudy (nuccioreggie):

Mark simplified three over five divided by five over eight.; his work is shown below. Identify where he made his error. Original division problem: three over five divided by five over eight. Step 1: three over five multiplied by eight over five. Step 2: three times eight over five times five. Step 3: twenty-four over five. Step 4: four and four over five.

OpenStudy (texaschic101):

(3/5) / (5/8) = (3/5) * (8/5) = 24/25 --- there is the mistake

OpenStudy (texaschic101):

he multiplied the numerators correctly but not the denominators

OpenStudy (texaschic101):

understand ? Because when dividing fractions, you " flip " what you are dividing by, then just multiply

OpenStudy (moazzam07):

so its step 3 right? @texaschic101